Goodman and Associates, a law firm, employs ABC. The following budgeted data for each of the...

  1. Goodman and Associates, a law firm, employs ABC. The following budgeted data for each of the activity cost pools is provided for the year 2019. (19 pts)

                                                         Estimated          Expected Use of

        Activity Cost Pools                Overhead   Cost Drivers per Activity

        Researching legal issues        $     31,500               900 research hours

        Preparing legal documents          480,000          30,000 pages

        Meeting with clients                1,760,000          8,800 professional hours

During 2019 the firm worked 660 research hours, 10,000 professional hours, and prepared 25,000 document pages.

Instructions

Compute the total overhead applied during 2019.
Answer #1

Answer:

The total overhead applied during 2019 is $2,423,100

Calculations:

Unit cost:

Estimated
Overhead
÷ Cost
Diver
= Unit cost
Researching legal issues   $31,500 ÷ 900 Research hours = $35 per research hour
Preparing legal documents $480,000 ÷ 30,000 Pages = $16 per pages
Meeting with clients $1,760,000 ÷ 8,800 Professional hours = $200 per professional hour

Overhead applied:

Actual x Unit cost = Applied overhead
Researching legal issues   660 Research hours x $35 per research hour = $23,100
Preparing legal documents 25,000 Pages x $16 per pages = $400,000
Meeting with clients 10,000 Professional hours x $200 per professional hour = $2,000,000
Total $2,423,100